Suppose a space probe moves away from the Earth at a speed 0.350c. It sends a radio wave message back to the Earth at a frequency of 1.50 GHz. At what frequency is the message received on the Earth?

1
Expert's answer
2022-05-30T15:18:37-0400
"f=f_0\frac{c}{c+v}\\
=1.50*10^9*\frac{c}{c+0.350c}=1.11*10^9\:\rm Hz=1.11\: GHz"